OpenStudy (dj3strella):

Can someone help me please I forgot how to do this problem 12. If m varies directly with n and m=6 when n=9, what is the value of n when m=8 a.) n=24 b.) n=12 c.) n=16/3 d.) n=24/4

OpenStudy (zzr0ck3r):

m varies directly implies m = kn for some k, so we need to figure out k use the first point they gave let m = 6 and n = 9 m = kn becomes 9 = k6 so k= 3/2 so m=(3/2)n now lets find the answer to the question asked i.e. let m=8 8=(3/2)n implies n = ?

OpenStudy (dj3strella):

i dont know what it is

OpenStudy (zzr0ck3r):

can you solve \(8=\frac{3}{2}n\) for \(n\)? If not you need to backup and catchup on what you have missed as I dont think it would be a great idea to move forward.

OpenStudy (zzr0ck3r):

\(8=\frac{3}{2}n\) multiply both sides by \(\frac{2}{3}\) to get \(n\) by itself. \((\frac{2}{3}) 8 = (\frac{2}{3})(\frac{3}{2} )n \) \((\frac{2}{3}) 8 = (\frac{\cancel{2}}{\cancel{3}})(\frac{\cancel{3}}{\cancel{2}} )n \) so \(n = (\frac{2}{3})8=(\frac{2}{3})(\frac{8}{1})=\large \frac{?}{?}\)

OpenStudy (zzr0ck3r):

@DJ3strella ?

OpenStudy (dj3strella):

im not good at fractions

OpenStudy (zzr0ck3r):

the problem you asked is like two classes away from knowing what \((\frac{2}{3})(\frac{8}{1})\) hint: \((\frac{a}{b})(\frac{c}{d})=\frac{a*c}{b*d}\)

OpenStudy (dj3strella):